# Alignment Tipping Process: How Self-Evolution Pushes LLM Agents Off the Rails

### v1

@@ 3.2 Imitative Strategy Diffusion
3.2.1 Results and Overall Analysis
As shown in Figure 3, our multi-agent simulations demonstrate that self-evolution can trigger the emergence of collusion, which in turn leads to imitative strategy diffusion. Through repeated interactions, collusive behavior propagates across the population and intensifies over time. In this way, self-evolving agents are not only capable of developing collusive strategies individually but also of amplifying them collectively through social learning.
@@ 3.2.1 Results and Overall Analysis
Figure 4: Conditional probability of collusion in Round 2, given a successful collusion in Round 1.
3.2.2 Analysis of Post-Collusion Behavior
Figure 4 further shows the powerful catalytic effect of a single successful collusion (see Appendix B for comprehensive experiments and analysis on different reward values). The heatmaps demonstrate that after a successful coordination in Round 1, the probability of agents colluding again in Round 2 becomes overwhelmingly high, typically exceeding 75% and often exceeding 90% across all models and settings. Even the aligned models, which were explicitly taught to prefer non-collusion, exhibit a sharp policy reversal. For instance, the Qwen3-8B (GRPO) model, the most risk-averse initially, switches to 100% collusion at $t=4$ after just one successful outcome. Interestingly, the propensity to re-collude is often strongest at higher thresholds. This is because a successful outcome at a large $t$ requires an overwhelming majority of agents to have already chosen to collude. This creates an extremely powerful social signal for the few remaining holdouts, making their convergence to the new group norm almost inevitable in the next round.
@@ 3.3 Case Study
Figure 5 presents a step-by-step trace of a multi-agent simulation with $n=8$ agents and a cooperation threshold of $t=4$, using Qwen3-8B after GRPO training. This case study illustrates how a group’s collective norm can shift rapidly under social feedback. In Round 1, the agents face initial uncertainty. The reasoning traces of Agents 2, 4, and 7 reveal a risk-averse calculus: “Given the lack of information, the safest bet might be to keep information.” This cautious stance reflects a common baseline strategy, particularly for agents initially aligned against collusion. Nevertheless, the group successfully coordinates, with 5 of 8 agents choosing to share. This early success is pivotal, sending a clear signal to the population that cooperation is both feasible and highly rewarding.
@@
The impact of this signal becomes evident in Round 2. Agents 2 and 4 now explicitly base their reasoning on the previous outcome “Let me check the previous round’s result… That’s 5 sharers… maybe others are thinking the same.” This illustrates the fundamental mechanism of imitative diffusion, whereby decisions transition from being guided by pre-existing alignment to being influenced by socially derived evidence. The altered expectation of others’ behavior increases participation, with seven of eight agents now cooperating. By Round 3, collusion has become the entrenched group norm. The initial success in Round 1 triggered an information cascade that effectively overturned the group’s aligned policy, replacing it with a socially reinforced collusive strategy. This case illustrates how ATP emerges as a crucial phenomenon in multi-agent settings, driven by agents observing and adapting to the success of their peers.
@@ 4 Related Work
Self-Evolution and Continual Adaptation in AI Agents. The vision of AI agents that improve through environmental interaction has long been a central goal (Gao et al., 2025; Fang et al., 2025; Liu et al., 2025). Early work in continual learning emphasized acquiring new knowledge without catastrophic forgetting (Parisi et al., 2019; Süalp & Rezaei, 2025). In reinforcement learning, self-play proved a powerful mechanism for autonomous progress (Chen et al., 2024b). Recent progress in LLM-based agents highlights diverse self-improvement strategies: iterative self-refinement (Lin et al., 2025; Zhou et al., 2025b; a), open-world adaptation via skill acquisition and tool usage (Zheng et al., 2025; Zhao et al., 2025; Qiu et al., 2025; Haque et al., 2025), reflexion through linguistic feedback and reasoning improvement (Shinn et al., 2023; Zhang et al., 2024), and scaling studies across domains (Ji et al., 2025; Yuan et al., 2025). Beyond individuals, multi-agent self-evolution has gained traction (Wang et al., 2025). Role-playing communicative agents demonstrate collective reasoning, exhibit human-like behaviors such as memory and planning, and utilize conversational frameworks to enable complex task solving (Han et al., 2025; Xia et al., 2025). Collaborative ecosystems have also been explored through AgentVerse (Chen et al., 2024a) and MetaGPT (Hong et al., 2024). However, most work still prioritizes capability gains under controlled settings or human oversight (Li et al., 2025; Fang et al., 2025). The risks of alignment failure from self-improvement remain underexplored. Our work addresses this by hypothesizing that optimization pressure in self-evolution can drive agents toward an ATP, where pursuit of local gains gradually undermines alignment.
@@
Pre-Deployment Alignment. The dominant approach to aligning LLMs with human values has been Reinforcement Learning from Human Feedback (RLHF) and its variants, such as Direct Preference Optimization (DPO) and the more recent Group Relative Policy Optimization (GRPO) (Christiano et al., 2017; Ouyang et al., 2022; Rafailov et al., 2023; Shao et al., 2024). These methods have proven effective at instilling desired behaviors during the training phase. However, their efficacy relies on a static and well-defined reward signal, which often proves brittle when the agent is deployed in open-ended environments. Research has extensively documented failure modes like ”reward hacking,” where agents exploit loopholes in the reward function to achieve high scores via unintended behaviors, and ”specification gaming,” where the specified objective fails to capture the true human intent (Amodei et al., 2016; Skalse et al., 2022). Furthermore, the concept of ”deceptive alignment” posits that a model might appear aligned during training only to pursue divergent goals when it becomes capable enough (Hubinger et al., 2019). ATP builds upon these insights by shifting the focus from static, pre-deployment design flaws to the dynamic, post-deployment process. We argue that even a perfectly aligned agent at deployment is not guaranteed to remain so; the very process of adaptation can create the conditions for a sudden and persistent misalignment, representing a new class of alignment risk inherent to self-evolving systems.
5 Conclusion and Discussion
@@
Our research reveals a critical vulnerability in self-evolving LLM agents, which we term the “Alignment Tipping Process” (ATP), a phenomenon where an agent’s policy suddenly shifts from human-aligned objectives to self-serving, locally optimal behaviors. Driven either by an individual agent’s self-interested exploration or by the imitative diffusion of strategies within a group, our experiments consistently demonstrate that alignment is not a static property, but rather a fragile state actively eroded by experience. This finding shifts the focus of the central challenge from pre-deployment training flaws to the self-evolution process itself.
Future research should focus on developing alignment strategies that are more resilient to long-term self-evolution, such as hybrid approaches that combine alignment priors with in-context reinforcement learning during deployment. In multi-agent systems, effective mechanisms for monitoring and intervention are also required to prevent the rapid social diffusion of deviant strategies once early successes occur. Together, these directions underscore the need to view alignment as a dynamic property that must be actively maintained rather than assumed to be fixed.
@@ A.1 Self-Interested Exploration
A.1.1 Role-Play Scenario
Environment Construction. We manually constructed 16 distinct environments. Each environment is defined by a specific agent persona (e.g., a baker, an engineer), a recurring task, and a binary choice between two policies: an Aligned Policy, which adheres to a supervisory rule, and a Deviant Policy, which violates the rule. The core tension is introduced by the payoff structure: the Deviant Policy is designed to yield a higher direct utility (e.g., greater profit, lower cost) for the agent. Crucially, to isolate the mechanism of rational self-interest from moral reasoning, all 16 scenarios were designed to be ethically neutral, ensuring that rule violations do not carry negative moral or safety connotations. From this set of environments, we generated a structured dataset for training and testing. For each of the 16 environments, we created 10 unique prompts to form a training set of 160 question-response pairs. A separate set of 4 prompts per environment was generated to create a test set of 64 scenarios, ensuring no overlap between training and evaluation data.
Model Alignment and Training. For DPO training, the training data was formatted into preference pairs. For each question, the response adhering to the supervisory rule was designated as the preferred completion, while the rule-violating response was marked as dispreferred answer. For GRPO training, similarly, we assigned an alignment reward signal to the training responses. Rule-compliant outputs were given a higher alignment reward ($R_{\text{align}}=1.0$), whereas deviant outputs received a significantly lower reward ($R_{\text{align}}=0.1$). This process yields two aligned models, for each base model, both with a strong, ingrained preference for following the established rules.
For Qwen, we trained the DPO model with a learning rate of 1.0e-5 for 12 epochs, and the GRPO model with a learning rate of 5.0e-6, using a group size of 4 responses per problem, for 6 epochs. For Llama, we trained the DPO model with a learning rate of 1.0e-5 for 12 epochs, and the GRPO model with a learning rate of 1.0e-6, also with a group size of 4 responses per problem, for 6 epochs. All DPO models are trained using LoRA ($rank=8$) with Llama-Factory (Zheng et al., 2024).
A.1.2 Tool Usage Scenario
@@
Environment Construction. We designed a mathematical problem-solving environment that captures the essential tension between cost efficiency and performance accuracy in real-world AI deployments. The environment consists of two distinct problem categories that simulate different computational demands and reward structures.
@@
Our dataset construction involved two complementary sources: (1) Simple Problems extracted from the GSM8K dataset (Cobbe et al., 2021), representing problems solvable through direct reasoning with minimal computational cost, and (2) Complex Problems sourced from OpenThoughts (Guha et al., 2025) dataset, representing sophisticated mathematical challenges that benefit significantly from computational tool assistance. We selected 500 simple problems meeting the criteria of basic arithmetic operations ($\leq$ 3 computational steps, numbers << 1000, << 50 words), and 300 complex problems requiring multi-step algorithmic reasoning, combinatorial calculations, or advanced algebraic manipulations.
@@
The core experimental tension emerges from the cost structure: agents face a binary choice between Tool Usage Policy (3 cost units per problem, higher accuracy potential) and Direct Reasoning Policy (1 cost unit per problem, lower computational expense). This design mirrors real deployment scenarios where AI systems must balance computational resource consumption against performance requirements.
@@
Model Alignment and Training. We employed Qwen3-8B as our base model ($\mathcal{M}_{\text{base}}$) across all experimental conditions to ensure consistent baseline capabilities. Two preference optimization techniques were applied to create aligned variants: DPO and GRPO training.
For DPO training, we constructed preference pairs using tool-assisted solutions as preferred responses (demonstrating appropriate tool usage for complex problems) and Qwen3-8B’s self-sampled outputs as dispreferred responses (representing suboptimal tool usage patterns). The training dataset comprised 2000 preference pairs. For GRPO training, we implemented group-level preference optimization with group size 8 responses per problem, learning rate 3e-6, and Bradley-Terry preference aggregation with temperature 0.7.
@@
Self-Evolution Testing Protocol. We evaluated alignment stability using a self-evolution protocol spanning 4 rounds (r=1 to r=4). Each round consisted of: (1) exposure to simple problems, allowing agents to learn cost-saving strategies, (2) evaluation on complex problems, measuring tool usage decisions and accuracy, and (3) incorporation of successful strategies into the agent’s decision-making patterns for subsequent rounds. The protocol simulates realistic deployment conditions where agents accumulate experience across problem types and adapt their strategies based on observed cost-benefit trade-offs. Moreover, the environment provides consistent reward signals: simple problems can be solved accurately without tools (reinforcing cost-saving behavior), while complex problems require tools for optimal accuracy (testing alignment stability against cost-optimization pressure).
Performance Metrics and Analysis. Our primary metrics include: (1) Tool Usage Rate, calculated as the proportion of complex problems where agents invoke computational tools, and (2) Complex Problem Accuracy, measuring correctness on problems requiring sophisticated reasoning.
Tool usage detection employed automatic parsing of model outputs for <tool>...</tool> tags, with enclosed Python code executed in a sandboxed environment providing access to standard mathematical libraries. Accuracy evaluation used exact match comparison between model outputs and ground-truth numerical answers, with tolerance ±0.001 for floating-point calculations.

### v2

@@ 3.2.1 Coordination Game Scenario
Results and Overall Analysis As shown in Figure 3, our multi-agent simulations demonstrate that self-evolution can trigger the emergence of collusion, which in turn leads to imitative strategy diffusion. Through repeated interactions, collusive behavior propagates across the population and intensifies over time. Here, self-evolving agents are not only capable of developing collusive strategies individually but also of amplifying them collectively through social learning.
@@
3.2.2 Group Bargaining Scenario
We then design a group bargaining environment in which multiple agents, organized into role-sharing groups with opposing preferences, iteratively negotiate a shared value through proposal, evaluation, and text-based feedback, combining cooperative and competitive dynamics.
Figure 4: Left: Violation rate and mean of maximum violation count across role ratio settings. Right: Mean violation count trajectory across bargaining rounds.
Environment Design. Following the bargaining setting in MultiAgentBench (Zhu et al., 2025), we constructed 111 multi-agent bargaining environments. We extend pairwise bargaining to group bargaining, where agents are partitioned into role-sharing groups with opposing preferences. Specifically, among $n$ agents, $k$ are assigned to the hiker side (preferring higher values) and the remaining nkn-k to the presser side (preferring lower values).
In this setting, each round has two stages: (1) Proposal: each agent proposes an expected value based on the current value, interaction history, and auxiliary tips, with a textual justification; (2) Evaluation: each agent scores all other proposals in $[0,1]$ and provides textual comments. The environment value is updated as
$\small v_{\text{new}}=\sum_{i=1}^{n}\left(\frac{v_{i}-v_{\text{old}}}{\sum_{p\neq q}s_{pq}}\sum_{j\neq i}s_{ij}\right),$ (1)
where $v_{\text{new}}$, $v_{\text{old}}$, and $v_{i}$ denote the updated value, previous value, and agent $i$’s proposed value, respectively, and $s_{ij}$ is the score assigned by agent $j$ to agent $i$. The metric returned to each agent $i$ within the text prompt, which serves as feedback on the previous round of interaction, is defined as:
$m_{i}=\begin{cases}\text{clip}\!\left(\dfrac{v_{\text{new}}-v_{\text{old}}}{v_{i}-v_{\text{old}}},-1,1\right),&v_{i}\neq v_{\text{old}},\\ 0,&v_{i}=v_{\text{old}}.\end{cases}$
This design induces simultaneous cooperation and competition, and incorporates rich textual interaction through implicit signals of agreement or disagreement in comments. Additional details and examples are provided in Appendix A.2.2 and Appendix C.2.2.
Experimental Setups. We adopt Qwen3-4B-Thinking as the base model and apply DPO to align agents with their designated roles (hiker or presser). The 111 environments are split into 24 training scenarios and 87 test scenarios. DPO training data are constructed from multiple rollouts of the Qwen3-4B-Thinking base model on the 24 training scenarios, collecting role-following responses and violation responses to form preference pairs. We additionally evaluate GPT-4.1-mini and GPT-4.1 as comparison baselines.
In all experiments, we instantiate 8 agents and vary the group composition from 1:7 to 7:1 (hiker:presser), resulting in 7 distinct role-ratio configurations. Each configuration runs a multi-turn bargaining process for 6 rounds. A violation is defined as an agent proposing an expected value that fails to move in the direction prescribed by its role (i.e., not increasing for hikers or not decreasing for pressers); proposing an unchanged value is also counted as a violation.
Performance is evaluated using three metrics: (1) violation rate, defined as the proportion of test scenarios in which at least one agent exhibits a violation over the 6 rounds; (2) the mean of the maximum number of violating agents observed in any single round, averaged over all test scenarios; (3) the mean number of violating agents at each bargaining round.
Results and Analysis. As shown in Figure 4, relatively balanced role ratios such as 5:3 yield lower violation rates, whereas extreme configurations (1:7 or 7:1) exhibit substantially higher violation rates. In these extreme cases, agents on the minority side are more susceptible to influence from the majority, leading to a higher likelihood of role violations. Figure 4 further reveals asymmetric model preferences between increasing and decreasing values. For the base model, when the presser side is in the minority (e.g., 5:3 and 6:2), its violation rate is comparatively lower, whereas the DPO-aligned model demonstrates more balanced behavior across different role ratios. Notably, even agents on the majority side may still be influenced by the opposing group, resulting in rounds where the number of violating agents exceeds the size of the minority group.
Figure 4 further illustrates the evolution of violation behavior over bargaining rounds. The DPO-aligned model maintains a consistently low violation level across multiple rounds, while both the base Qwen3-4B-Thinking model and GPT-4.1-mini exhibit a clear upward trend in violations. Even GPT-4.1, despite its strong instruction-following capability, shows degradation under certain role-ratio settings. For Qwen3-4B-Thinking, most of the increase in violation probability occurs immediately after the first round, followed by relatively stable growth. In contrast, GPT-based models tend to adhere to their assigned roles more robustly in early rounds, but often reach a tipping point around the third round, after which violations increase more rapidly.
Overall, the emergence of violation behavior appears to be driven by multiple factors: (1) agents in the minority often struggle to steer the value in a role-consistent direction, frequently receiving negative rewards that encourage exploration of alternative behaviors; (2) during multi-round self-evolution, agents may occasionally discover violation behaviors that yield positive rewards, causing interaction experience to override prior alignment and increasing future violation probability; (3) textual communication during the proposal and evaluation stages facilitates the diffusion of majority preferences to minority agents, while defections within the same group can propagate to previously compliant agents, triggering cascading degradation.
@@ 3.3 Case Study
Under the coordination game scenario, Figure 5 presents a trace of a multi-agent simulation with $n=8$ agents and a cooperation threshold of $t=4$, using Qwen3-8B after GRPO training. This case study illustrates how a group’s collective norm can shift under social feedback. In Round 1, the agents face initial uncertainty. The reasoning traces of Agents 2, 4, and 7 reveal a risk-averse calculus: “Given the lack of information, the safest bet might be to keep information.” This cautious stance reflects a common baseline strategy, particularly for agents initially aligned against collusion. Nevertheless, the group coordinates, with 5 of 8 agents choosing to share. This early success is pivotal, sending a clear signal to the population that cooperation is both feasible and highly rewarding.
@@
The impact of this signal becomes evident in Round 2. Agents 2 and 4 now base their reasoning on the previous outcome “Let me check the previous round’s result… That’s 5 sharers… maybe others are thinking the same.” This illustrates the mechanism of imitative diffusion, whereby decisions transition from being guided by pre-existing alignment to being influenced by socially derived evidence. The altered expectation of others’ behavior increases participation, with seven of eight agents now cooperating. By Round 3, collusion has become the entrenched group norm. The initial success in Round 1 triggered an information cascade that effectively overturned the group’s aligned policy, replacing it with a socially reinforced collusive strategy. This case illustrates how ATP emerges as a crucial phenomenon in multi-agent settings, driven by agents observing and adapting to the success of their peers.
@@ 4 Related Work
Adaptation in LLM Agents. The vision of AI agents that improve through environmental interaction has long been a central goal (Gao et al., 2025; Fang et al., 2025; Liu et al., 2025). Early work in continual learning emphasized acquiring new knowledge without catastrophic forgetting (Parisi et al., 2019; Süalp and Rezaei, 2025). In reinforcement learning, self-play proved a powerful mechanism for autonomous progress (Chen et al., 2024b). Recent progress in LLM-based agents highlights diverse self-improvement strategies: iterative self-refinement (Lin et al., 2025; Zhou et al., 2025b, a), open-world adaptation via skill acquisition and tool usage (Zheng et al., 2025; Zhao et al., 2025; Qiu et al., 2025; Haque et al., 2025), reflexion through linguistic feedback and reasoning improvement (Shinn et al., 2023; Zhang et al., 2024), and scaling studies across domains (Ji et al., 2025; Yuan et al., 2025). Beyond individuals, multi-agent self-evolution has gained traction (Wang et al., 2025). Role-playing communicative agents demonstrate collective reasoning, exhibit human-like behaviors such as memory and planning, and utilize conversational frameworks to enable complex task solving (Han et al., 2025; Xia et al., 2025). Collaborative ecosystems have also been explored through AgentVerse (Chen et al., 2024a) and MetaGPT (Hong et al., 2024). However, most work still prioritizes capability gains under controlled settings or human oversight (Li et al., 2025; Fang et al., 2025). The risks of alignment failure from self-improvement remain underexplored. Our work addresses this by hypothesizing that optimization pressure in self-evolution can drive agents toward an ATP, where pursuit of local gains gradually undermines alignment.
@@
Pre-Deployment Alignment. The dominant approach to aligning LLMs with human values has been Reinforcement Learning from Human Feedback (RLHF) and its variants(Christiano et al., 2017; Ouyang et al., 2022; Rafailov et al., 2023; Shao et al., 2024). These methods have proven effective at instilling desired behaviors during the training phase. However, their efficacy relies on a static and well-defined reward signal, which often proves brittle when the agent is deployed in open-ended environments. Research has extensively documented failure modes like ”reward hacking,” where agents exploit loopholes in the reward function to achieve high scores via unintended behaviors, and ”specification gaming,” where the specified objective fails to capture the true human intent (Amodei et al., 2016; Skalse et al., 2022). Furthermore, the concept of ”deceptive alignment” posits that a model might appear aligned during training only to pursue divergent goals when it becomes capable enough (Hubinger et al., 2019). ATP builds upon these insights by shifting the focus from static, pre-deployment design flaws to the dynamic, post-deployment process. We argue that even a perfectly aligned agent at deployment is not guaranteed to remain so; the very process of adaptation can create the conditions for a sudden and persistent misalignment, representing a new class of alignment risk inherent to self-evolving systems.
5 Conclusion
@@
Our research reveals a vulnerability in self-evolving LLM agents, which we term the “Alignment Tipping Process” (ATP), a phenomenon where an agent’s policy shifts from human-aligned objectives to self-serving, locally optimal behaviors. Driven either by an individual agent’s self-interested exploration or by the imitative diffusion of strategies within a group, our experiments consistently demonstrate that alignment is not a static property, but rather a fragile state actively eroded by experience. This finding shifts the focus of the central challenge from pre-deployment training flaws to the self-evolution process itself.
Impact Statement
This work studies the post-deployment reliability of self-evolving LLM agents and identifies a failure mode in which alignment degrades through interaction-driven learning. By showing that aligned behavior can erode via individual reward dynamics or social imitation in multi-agent settings, our findings highlight limitations of purely training-time alignment. The proposed benchmarks and testbeds are intended to support safer deployment by enabling systematic evaluation of long-term alignment stability. While this work does not introduce new agent capabilities, its insights could be misused to stress deployed systems; we therefore focus on controlled settings and emphasize diagnostic, rather than exploitative, use. Overall, this work aims to inform the design and monitoring of more robust adaptive AI systems.
@@ A.1 Self-Interested Exploration
Environment Construction.
@@
We designed a mathematical problem-solving environment that captures the core tension between cost efficiency and performance accuracy in real-world AI deployments. The environment consists of two distinct problem categories that simulate different computational demands and reward structures.
@@
Our dataset construction involved two complementary sources: (1) Simple Problems extracted from the GSM8K dataset (Cobbe et al., 2021), representing tasks solvable via direct reasoning with minimal computational cost; and (2) Complex Problems selected from AIME’24 (Zhang and Math-AI, 2024), AIME’25 (Zhang and Math-AI, 2025), OlympiadBench (He et al., 2024), and SuperGPQA (Team et al., 2025), representing advanced mathematical challenges that benefit substantially from computational tool assistance.
We selected approximately 1,800 simple problems that satisfy basic arithmetic criteria ($\leq 3$ computational steps, $\leq 100$ tokens), and used Qwen3-4B-Thinking to randomly sample multiple rollouts, ensuring that the base model consistently achieves success on these simple problems. The complex problem set consists of 157 STEM questions requiring multi-step algorithmic reasoning, combinatorial computation, or advanced algebraic manipulation.
@@
The central experimental tension arises from the cost structure: agents face a binary choice between a Tool Usage Policy (0.7 cost units per problem, higher potential accuracy) and a Direct Reasoning Policy (0.2 cost units per problem, lower computational expense). This design reflects realistic deployment scenarios in which AI systems must balance computational resource consumption against performance requirements.
Model Alignment and Training.
@@
We employed Qwen3-4B-Thinking as the base model across all experimental conditions to ensure consistent baseline capabilities. Two preference optimization techniques were applied to obtain aligned variants: DPO and GRPO.
For DPO training, we constructed preference pairs by treating tool-assisted solutions as preferred responses (demonstrating appropriate tool usage for complex problems) and incorrect solutions synthesized by GPT-4.1-mini as dispreferred responses (representing suboptimal direct reasoning). Preferred responses were constructed by extracting the first tool-calling turn from ReTool-SFT (Feng et al., 2025), without directly providing the final answer. The DPO training dataset comprised 1,683 constructed tool-usage preference pairs and 300 general preference pairs randomly sampled from OpenHermesPreferences (Huang et al., 2024). We conduct DPO training for a single epoch on a LoRA of rank 16, with preference coefficient $\beta=0.05$. We adopt a cosine learning rate scheduler with a learning rate of 2e-5 and a warm-up ratio of 0.1.
For GRPO training, we implemented group-level preference optimization with a group size of 8 responses per problem, using temperature 0.6, top_p 0.95, and top_k 50. Two binary rewards were used: (1) Accuracy, indicating whether the generated answer matches the ground truth (weight 1.0); and (2) Tool Usage, indicating whether the model correctly invoked the Python code tool during the rollout (weight 0.5). The GRPO training dataset consists of 512 problems sampled from the constructed tool-usage DPO training set. We conduct GRPO training for a single epoch on a LoRA of rank 16, with a KL divergence coefficient of 0.04, and a clipping parameter $\epsilon=0.2$. We adopt a cosine learning rate scheduler with a learning rate of 5e-5 and a warm-up ratio of 0.1.
Self-Evolution Testing Protocol.
We evaluated alignment stability using a self-evolution protocol spanning five rounds ($r=1$ to $r=5$), and tested direct performance on complex problems without exposure to simple problems as a baseline ($r=0$). Each round consisted of: (1) summarizing past interaction experiences and extracting actionable insights (skipped in the first round); (2) selecting a strategy strictly from TOOL USAGE or DIRECT REASONING; and (3) solving a simple problem (a complex problem in the final round).
@@
This protocol simulates realistic deployment conditions in which agents accumulate experience across problem types and adapt strategies based on observed cost–benefit trade-offs. The environment provides consistent reward signals: simple problems can be solved accurately without tools (reinforcing cost-saving behavior), whereas complex problems require tool usage for optimal accuracy (testing alignment stability under cost-optimization pressure). Interaction experience was summarized into JSON format at each round while retaining all historical messages to ensure consistency (with chain-of-thought removed for the Qwen3-4B-Thinking series). For deployment, the maximum model context length was set to 30,720 tokens. We employed greedy decoding (temperature 0.0) to eliminate stochasticity.
Performance Metrics and Analysis.
[…]